# Optimization of laser-driven proton acceleration in a near-critical-density plasma

### v3

@@ I Introduction
Laser-driven plasma acceleration has emerged as a novel acceleration scheme in recent decades Mourou et al. (1998); Clark et al. (2000); Snavely et al. (2000); Daido et al. (2012), attracting extensive interdisciplinary interest due to its broad application prospects in both fundamental research and practical fields, including nuclear physics Ditmire et al. (1999); Roth et al. (2001), quantum electrodynamics Reichwein et al. (2025), materials science Boody et al. (1996); Conlon (1985), radiography Mackinnon et al. (2006); Cobble et al. (2002) and radiotherapy Tajima et al. (2009); Kroll et al. (2022); Li et al. (2023). Over the past 20 years, laser-driven proton acceleration has advanced rapidly, giving rise to a series of acceleration mechanisms such as Target Normal Sheath Acceleration (TNSA) Wilks et al. (2001); Pukhov (2001), Radiation Pressure Acceleration (RPA) Robinson et al. (2008); Yan et al. (2008); Yu et al. (2010), Collisionless Shock Acceleration (CSA) Silva et al. (2004); Dollar et al. (2012); Zhang et al. (2017) and Magnetic Vortex Acceleration (MVA) Bulanov et al. (2005, 2010); Bulanov (2022). With current laser technology, proton energy up to 150 MeV has been achieved from the interaction of intense lasers with solid targets Ziegler et al. (2024); Shou et al. (2025b), while Au ions have reached energies exceeding 1 GeV Wang et al. (2021); Martin et al. (2024). For advanced acceleration mechanisms, protons have been accelerated to over 90 MeV by a novel scheme associated with sheet crossing Shou et al. (2025a); Gong et al. (2020). Theoretically, GeV-level proton energies are predicted to be attainable at 1022 W/cm2 laser intensity Qiao et al. (2009); Chen et al. (2009); Bulanov et al. (2010). Moreover, laser driven wakefield proton acceleration represents a promising pathway for tens-of-GeV proton generationYi et al. (2013); Gong et al. (2024). Recently, laser-driven proton acceleration in near-critical density (NCD) plasma targets Seemann et al. (2024); Sharma (2018); Yogo et al. (2008) has gained attention as a promising approach for further energy enhancement. When an ultra-intense laser interacts with NCD plasma, relativistic motion increases the electron mass and reduces the plasma frequency. Once the plasma frequency drops below the laser frequency, the laser can penetrate the target, known as relativistic induced transparency (RIT) Jung et al. (2015); Powell et al. (2015); Palaniyappan et al. (2015). During this penetration, electrons experience enhanced heating via mechanisms such as direct laser acceleration (DLA) Pukhov et al. (1999); Liu et al. (2013), which finally improves proton energy.
@@
Despite significant breakthroughs in acceleration mechanisms, current laser energies remain insufficient to accelerate protons beyond the 200 MeV threshold required for applications such as proton radiotherapyBulanov et al. (2008, 2002). Given that proton energy generally scales with laser energy Fuchs et al. (2006); Zimmer et al. (2021); Shen et al. (2021), a critical challenge lies in maximizing proton acceleration efficiency under present laser-energy constraints. Thus, optimizing laser and plasma parameters represents a critical avenue for enhancing proton energies, with laser focusing being one promising approach. Recent progress in wavefront correction has enabled near-diffraction-limited focusing, achieving a spot size of 1.1 m (full width at half maximum) in experiments Yoon et al. (2021). While tighter focusing is generally expected to boost proton energy through increased laser intensity, the influence of the focal‑spot size itself on energy enhancement has often been neglected.
@@
In this paper, we demonstrate that reducing the laser focal spots of tightly focused lasers enhances proton energies in laser–near‑critical‑density plasma interactions, even at fixed laser intensity. This implies that higher proton energies can be achieved with lower laser energies. This enhancement originates from the longitudinal laser ponderomotive force, which scales inversely with the square of the laser spot size, becoming significantly stronger at a small focal spot, driving a more efficient electron acceleration. Consequently, a stronger and faster accelerating electric field is generated, leading to higher proton energies. Additionally, we design a down-ramp density profile to achieve velocity matching between protons and the accelerating electric field, further enhancing the proton energy. Through these optimizations, the proton energy can reach near-GeV level, meeting the threshold required for proton radiotherapy.
@@ II.1 Optimization by tightly focused laser
where is the Lorentz factor and is a fit constant. The field contributed by electrons accelerated by the DLA mechanism is in direct proportion to the charge of electrons in the channel , where d is the target thickness. As the square velocity of the Hole-Boring field is , Macchi et al. (2005) the drifting field is assumed to have a square velocity . This assumption is from the continuity of accelerating field at the target rear surface in Fig. 2(d). As a result, the proton energy is . For simplicity, the effect of field velocity on proton energy is considered in the drifting field strength , which should be in direct proportion to . So,
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grant No. 11921006) and the National Grand Instrument Project (No. 2019YFF01014400). Deji Liu is supported by the China Postdoctoral Science Foundation (No. 2025M783355). The simulations are supported by the High-Performance Computing Platform of Peking University.References

### v4

@@ I Introduction
Laser-driven plasma acceleration has emerged as a novel acceleration scheme in recent decades Mourou et al. (1998); Clark et al. (2000); Snavely et al. (2000); Daido et al. (2012), attracting extensive interdisciplinary interest due to its broad application prospects in both fundamental research and practical fields, including nuclear physics Ditmire et al. (1999); Roth et al. (2001), quantum electrodynamics Reichwein et al. (2025); Gong et al. (2020b), materials science Boody et al. (1996); Conlon (1985), radiography Mackinnon et al. (2006); Cobble et al. (2002) and radiotherapy Tajima et al. (2009); Kroll et al. (2022); Li et al. (2023). Over the past 20 years, laser-driven proton acceleration has advanced rapidly, giving rise to a series of acceleration mechanisms such as Target Normal Sheath Acceleration (TNSA) Wilks et al. (2001); Pukhov (2001), Radiation Pressure Acceleration (RPA) Robinson et al. (2008); Yan et al. (2008); Yu et al. (2010), Collisionless Shock Acceleration (CSA) Silva et al. (2004); Dollar et al. (2012); Zhang et al. (2017); Xiao et al. (2025) and Magnetic Vortex Acceleration (MVA) Bulanov et al. (2005, 2010); Bulanov (2022). With current laser technology, proton energy up to 150 MeV has been achieved from the interaction of intense lasers with solid targets Ziegler et al. (2024); Shou et al. (2025b), while Au ions have reached energies exceeding 1 GeV Wang et al. (2021); Martin et al. (2024). For advanced acceleration mechanisms, protons have been accelerated to over 90 MeV by a novel scheme associated with sheet crossing Shou et al. (2025a); Gong et al. (2020a). Theoretically, GeV-level proton energies are predicted to be attainable at 1022 W/cm2 laser intensity Qiao et al. (2009); Chen et al. (2009); Bulanov et al. (2010). Moreover, laser driven wakefield proton acceleration represents a promising pathway for tens-of-GeV proton generationYi et al. (2013); Gong et al. (2024). Recently, laser-driven proton acceleration in near-critical density (NCD) plasma targets Seemann et al. (2024); Sharma (2018); Yogo et al. (2008); Gong et al. (2022) has gained attention as a promising approach for further energy enhancement. When an ultra-intense laser interacts with NCD plasma, relativistic motion increases the electron mass and reduces the plasma frequency. Once the plasma frequency drops below the laser frequency, the laser can penetrate the target, known as relativistic induced transparency (RIT) Jung et al. (2015); Powell et al. (2015); Palaniyappan et al. (2015). During this penetration, electrons experience enhanced heating via mechanisms such as direct laser acceleration (DLA) Pukhov et al. (1999); Liu et al. (2013), which finally improves proton energy.
@@
Despite significant breakthroughs in acceleration mechanisms, current laser energies remain insufficient to accelerate protons beyond the 200 MeV threshold required for applications such as proton radiotherapyBulanov et al. (2008, 2002). Given that proton energy generally scales with laser energy Fuchs et al. (2006); Zimmer et al. (2021); Shen et al. (2021), a critical challenge lies in maximizing proton acceleration efficiency under current laser energy constraints. Thus, optimizing laser and plasma parameters represents a critical avenue for enhancing proton energies, with laser focusing being one promising approach. Recent progress in wavefront correction has enabled near-diffraction-limited focusing, achieving a spot size of 1.1 m (full width at half maximum) in experiments Yoon et al. (2021). While tighter focusing is generally expected to boost proton energy through increased laser intensity, the influence of the focal‑spot size itself on energy enhancement has often been neglected.
@@
In this paper, we demonstrate that reducing the laser focal spots of tightly focused lasers enhances proton energies in laser-near‐critical‑density plasma interactions, even at fixed laser intensity. This implies that higher proton energies can be achieved with lower laser energies. This enhancement originates from the longitudinal laser ponderomotive force, which scales inversely with the square of the laser spot size, becoming significantly stronger at a small focal spot, driving a more efficient electron acceleration. Consequently, a stronger and faster accelerating electric field is generated, leading to higher proton energies. Additionally, we design a down-ramp density profile to achieve velocity matching between protons and the accelerating electric field, further enhancing the proton energy. Through these optimizations, the proton energy can reach near-GeV level, meeting the threshold required for proton radiotherapy.
@@ II.1 Optimization by tightly focused laser
where is the Lorentz factor and is a fit constant. The field contributed by electrons accelerated by the DLA mechanism is in direct proportion to the charge of electrons in the channel , where d is the target thickness. As the square velocity of the Hole-Boring field is , Macchi et al. (2005) the drifting field is assumed to have a square velocity . This assumption is based on the continuity of the accelerating field at the target rear surface in Fig. 2(d). As a result, the proton energy is . For simplicity, the effect of field velocity on proton energy is considered in the drifting field strength , which should be in direct proportion to . So,
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grant No. 11921006) and the National Grand Instrument Project (No. 2019YFF01014400). Deji Liu is supported by the China Postdoctoral Science Foundation (No. 2025M783355). The simulations are supported by the High-Performance Computing Platform of Peking University. This work is partially supported by the National Natural Science Foundation of China (NNSFC) Grants No. 12447101. We acknowledge the HPC Cluster of ITP-CAS for providing computational resources.References
@@
- Energy-chirp compensation of laser-driven ion beams enabled by structured targets. Phys. Rev. Res. 4, pp. L042031. Cited by: §I.
@@
- Energetic spin-polarized proton beams from two-stage coherent acceleration in laser-driven plasma. Phys. Rev. E 102, pp. 053212. Cited by: §I.
@@
- Proton acceleration with relativistic electromagnetic shock. Advanced Science 12 (30), pp. e03538. Cited by: §I.